(a) Find the probability that the company will meet ts goal on a particular Probablaty Probability mm on a
Answer #1

a)

expected number of failure in 100 miles =6*2=12 =\lambda

hence P(meet its goal)=P(X<=6) =\sum_{x=0}^{6}e^{-\lambda }\lambda ^{x}/x! =0.0458

b)P(not meet its goal) =1-P(meet goal) =1-0.0458 =0.9542

c)

expected number of failure in 200 miles =6*4 =24=\lambda

P(X<=6) =\sum_{x=0}^{6}e^{-\lambda }\lambda ^{x}/x! =0.0000

d)

expected number of failure in 150 miles =6*3=18 =\lambda

P(X>12) =1-P(X<=12) =1-\sum_{x=0}^{12}e^{-\lambda }\lambda ^{x}/x! =1-0.0917 =0.9083